OpenStudy (anonymous):

According to a USA Today Gallup poll, there is a racial divide on views of the Trayvon Martin/George Zimmerman case in Sanford, FL. The poll showed that 73% of blacks polled believe George Zimmerman would have been arrested if Trayvon Martin was white, while 771 of 2,334 whites polled agree. What percent of whites polled agree that George Zimmerman would have been arrested?

OpenStudy (anonymous):

To get a percentage, you take the number agreed (in this case 771) and divide by the total number (2334) then multiply by 100 because you will get a decimal. (771/2334) * 100 = answer in %

OpenStudy (anonymous):

33% ?

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

wooooooo thanks

OpenStudy (anonymous):

can you please help me with one more question

OpenStudy (anonymous):

Frederick needs to borrow $3000 for a down payment on a new car. His uncle decided to loan him the money as a 3.5% simple interest loan. Frederick doesn’t want to pay more than $210 in interest. How many years does Frederick have to pay back the loan with interest?

OpenStudy (anonymous):

Okay, the equation for a simple interest loan is I = PRT where I is the total interest, P is the principle, R is the rate (in decimal form so divide the 3.5 by 100) and T is the time in years. In this case, I = 210, P = 3000, R = 0.035 and we are solving for T.

OpenStudy (anonymous):

210=3000*0.035*T?

OpenStudy (anonymous):

is it 105

OpenStudy (anonymous):

You have the equation right, but the answer shouldn't be that high.

OpenStudy (anonymous):

Oh, the 105 you have is the 3000 * 0.035. So now you are at 210 = 105*T